### v1

# NutePrune: Efficient Progressive Pruning with Numerous Teachers for Large Language Models

Abstract
The considerable size of Large Language Models (LLMs) presents notable deployment challenges, particularly on resource-constrained hardware. Structured pruning, offers an effective means to compress LLMs, thereby reducing storage costs and enhancing inference speed for more efficient utilization. In this work, we study data-efficient and resource-efficient structure pruning methods to obtain smaller yet still powerful models. Knowledge Distillation is well-suited for pruning, as the intact model can serve as an excellent teacher for pruned students. However, it becomes challenging in the context of LLMs due to memory constraints. To address this, we propose an efficient progressive Numerous-teacher pruning method (NutePrune). NutePrune mitigates excessive memory costs by loading only one intact model and integrating it with various masks and LoRA modules, enabling it to seamlessly switch between teacher and student roles. This approach allows us to leverage numerous teachers with varying capacities to progressively guide the pruned model, enhancing overall performance. Extensive experiments across various tasks demonstrate the effectiveness of NutePrune. In LLaMA-7B zero-shot experiments, NutePrune retains 97.17% of the performance of the original model at 20% sparsity and 95.07% at 25% sparsity.
NutePrune: Efficient Progressive Pruning with Numerous Teachers for Large Language Models
Shengrui Li1 Tsinghua University Microsoft Research Asia lsr22@mails.tsinghua.edu.cn                        Xueting Han2 Microsoft Research Asia chrihan@microsoft.com                        Jing Bai Microsoft Research Asia jbai@microsoft.com
11footnotetext: Work is done during internship at Microsoft Research Asia.22footnotetext: Corresponding author.
1 Introduction
Large Language Models (LLMs) excel in language tasks OpenAI (2023); Touvron et al. (2023); Thoppilan et al. (2022); Scao et al. (2022), but their substantial size poses deployment and inference challenges Frantar et al. (2022). Techniques like model pruning Molchanov et al. (2016), knowledge distillation Jiao et al. (2019), and quantization Dettmers et al. (2023) have been proposed to address computational demands. The exploration of LLM pruning, especially structured pruning Frantar and Alistarh (2023), holds great significance. Structured pruning reduces model size by removing coherent parameter groups, cutting inference costs on standard hardware. But it is more challenging than unstructured pruning in retaining the capabilities of LLMs Hoefler et al. (2021). Existing methods either adopt data-efficient approaches, causing a performance decline Ma et al. (2023), or require extensive post-training to recover model performance Xia et al. (2023). In this work, we investigate efficient methods to prune the model to higher sparsity without significant performance decline.
Figure 1: The advantage of our NutePrune. Left: Progressive distillation guides the student with teachers from easy to hard to avoid large capacity gap harming learning. But it suffers from multiple-fold costs of loading numerous teachers. Right: Our NutePrune leverages models with varying sparsity, enabling progressive distillation with negligible additional cost.
Knowledge distillation (KD) aims to train a more compact student model with supervision from a larger teacher model Sanh et al. (2019); Gou et al. (2021). It’s widely adopted and proven highly effective in the field of LLMs. Progressive learning, utilizing intermediate teachers with a reduced gap in capabilities, has been demonstrated to improve performance in KD Xiang et al. (2020). Previous work has shown that pruning with a distillation objective can improve performance Xia et al. (2022). Distillation is particularly suitable for pruning since the full original model inherently serves as an excellent teacher for the pruned model Sanh et al. (2020), which can offer a more detailed supervisory signal than conventional supervised training, enhancing the effectiveness of pruning with limited data Lagunas et al. (2021).
However, applying this method in the realm of LLMs proves challenging. Given the vastness of an LLM, loading it onto GPUs consumes a substantial amount of memory. Introducing an additional teacher model requires twice the memory, making it impractical with limited memory resources. Furthermore, relying on a single teacher may not be the best practice Liu et al. (2020); Wu et al. (2021). With the increasing gap of sparsity between teacher and student, the capacity gap is also widening, which toughens distillation. Employing multiple teachers with varying capacities can enhance the transfer of knowledge to students Yuan et al. (2021). However, when it comes to the distillation of LLMs, memory consumption of multiple teachers becomes an even more pressing concern.
In this paper, we address the above challenges with an efficient progressive Numerous-teacher pruning method (NutePrune). Our motivation is demonstrated in Figure 1. NutePrune aims to diminish the capacity gap between the full teacher model and the highly sparse student, thereby alleviating the difficulty of distillation Su et al. (2021); Mukherjee et al. (2023); Xiang et al. (2020). Instead of relying solely on a single full teacher, we instruct the student with many teachers with varying sparsity. To achieve this, we formulate pruning as a optimization problem where we learn masks to prune sub-modules while updating model parameters through LoRA Hu et al. (2021). Specially, we load an intact model, serving dual roles as both a teacher and a student. In teacher mode, we incorporate the original model with collected frozen low-sparsity masks and corresponding LoRA modules. And in student mode, we incorporate it with learnable high-sparsity masks and LoRA modules. Since the masks and LoRA modules are highly parameter efficient, we collect and leverage numerous modules with different sparsity to incorporate numerous teachers and progressively prune the student. Our contributions can be summarized as follows:
  • We propose a novel distillation method that progressively guide the student using numerous teachers with varying sparsity to narrow the capacity gap. Through progressive KD, we achieve higher model sparsity without significant performance decline on limited data.
  • Our NutePrune only loads one intact model and switch it between teacher and student modes by incorporating various masks and LoRA modules. This novel efficient distilling method for pruning enables using numerous teachers and introduces no extra memory cost, which is especially critical for LLMs.
  • Extensive experiments across perplexity metric, commonsense reasoning, and MMLU demonstrate the effectiveness of our method.
2 Related Works
Pruning for LLMs
For LLMs, SparseGPT Frantar and Alistarh (2023) and WANDA Sun et al. (2023) employ unstructured pruning methods, while N:M sparsity Zhou et al. (2021) is considered semi-structured. Despite the effectiveness of these methods, their intricate structures do not yield significant inference speedup on standard hardware Frantar and Alistarh (2023). Consequently, structured pruning has emerged as a recent consensus. CoFi Xia et al. (2022) and nn pruning Lagunas et al. (2021) are proposed for smaller language models like BERT Devlin et al. (2018), often designed for specific tasks. CoFi loads both the teacher and student models, which is impractical for LLMs. Sheared-LLaMA Xia et al. (2023) proposes pruning LLMs using a dynamic pre-training method, enhancing performance through extensive data and training resources.
However, concerns persist regarding limited memory and training resources for LLMs. In a pioneering effort, LLM-Pruner Ma et al. (2023) prunes LLMs in one-shot and utilizes LoRA Hu et al. (2021) for fine-tuning. LoRAPrune Zhang et al. (2023) employs iterative pruning, replacing gradients on full weights with gradients on LoRA to calculate group importance. Compresso Guo et al. (2023) leverages LoRA and elaborately designed prompts for training and inference. Meanwhile, LoRAShear Chen et al. (2023) employs LoRA and a dynamic fine-tuning scheme to recover knowledge.
Knowledge Distillation (KD) for LLMs
KD Hinton et al. (2015) has emerged as a vital technique to reduce inference costs while maintaining performance quality in the context of LLMs. Prior work of KD Taori et al. (2023); Fu et al. (2023) mostly focuse on black-box KD, using teacher’s generations to fine-tune the student. With the rise of open-source LLMs Zhang et al. (2022); Touvron et al. (2023), interest in white-box KD is growing. White-box KD, leveraging teacher weights and logits, provides richer supervision signals, enhancing language abilities Agarwal et al. (2023); Gu et al. (2023); Wen et al. (2023). Despite progress on small language models, significant performance gaps between large and small models persist Achiam et al. (2023); Anil et al. (2023).
Progressive knowledge distillation Xiang et al. (2020) has proven effective by using intermediate teachers to bridge the capacity gap with LLMs, especially in scenarios reliant on data generated by multiple teachers Mukherjee et al. (2023). Orca Mukherjee et al. (2023) first learns from easier examples from ChatGPT and then from harder ones from GPT-4, enhancing performance for smaller students in KD. However, applying white-box KD to LLMs poses challenges due to substantial memory requirements for loading both teacher and student models. This challenge becomes even more difficult when attempting to load multiple teachers.
3 Methodology
Figure 2: The overall framework of NutePrune. The pruned model is frozen and incorporated with learnable masks and LoRA. During pruning, the model is guided by numerous teachers. Before pruned to the target sparsity (e.g. 30%), it learns from teachers with a fixed capacity gap. Once the target sparsity is achieved, it continues to learn from all previous teachers from weak to strong. All these teachers are derived from snapshots of the student model itself. Since only the mask and LoRA modules are snapshotted, the additional memory cost is negligible.
In this section, we first introduce how our NutePrune enables efficient knowledge distillation for structured pruning in 3.1. Then, to narrow capacity gap during distillation, we introduce the progressive knowledge distillation method that collects and incorporates numerous teachers in 3.2. The overview framework is illustrated in Figure 2.
3.1 Efficient Distillation for Structured Pruning
We formulate structure pruning as a constrained optimization problem where we optimize masks and model to prune the structure to a target sparsity while maximizing performance. We use distillation loss instead of original language model loss and use $L_{0}$ regularization to control the pruned sparsity. To mitigate substantial memory consumption during LLM training, we utilize LoRA for model updates, making pruning the process of training these masks and LoRA parameters.
Learning masks to control the pruned structure
We allow for three types of structured pruning: attention heads, FFN intermediate dimensions, and hidden dimensions. It is achieved by learning corresponding masks $\mathbf{z}_{head},\mathbf{z}_{int},\mathbf{z}_{hid}\in\{0,1\}$. Formally, the multi-head attention module $\operatorname{MHA}(x)$ and feed-forward networks $\operatorname{FFN}(x)$ of layer $l$ are pruned as:
$\operatorname{MHA}^{l}(X)=\mathbf{z}_{hid}\cdot\sum_{h=1}^{N_{head}}\mathbf{z}_{head}^{l,h}\operatorname{Att}^{l,h}(X).$ (1)
$\begin{split}\operatorname{FFN}^{l}&(X)=\mathbf{z}_{hid}\\ &\cdot W_{D}^{l}\left(\mathbf{z}_{int}^{l}\cdot W_{U}^{l}(X)\cdot\operatorname{Act}(W_{G}^{l}(X))\right)\end{split}$ (2)
where $\operatorname{Att}()$ is the attention module and $\operatorname{Act}()$ is the activation. $W_{D},W_{U},W_{G}$ are down projection, up projection, and gating projection.
During mask training, we calculate the remaining size to obtain the expected sparsity $\hat{s}$:
$\begin{split}\hat{s}(\mathbf{z})&=\frac{1}{M}\cdot 4\cdot d_{h}\cdot\sum_{l}^{L}\sum_{h}^{N_{head}}\sum_{k}^{d}\mathbf{z}_{head}^{l,h}\mathbf{z}_{hid}^{k}\\ &+\frac{1}{M}\cdot 3\cdot\sum_{l}^{L}\sum_{i}^{d_{int}}\sum_{k}^{d}\mathbf{z}_{int}^{l,i}\mathbf{z}_{hid}^{k},\end{split}$ (3)
where $M$ denotes full model size. $L$ is number of layers. $d_{h},N_{head},d,d_{int}$ are head dimension, number of head, hidden dimension, and intermediate dimension, correspondingly.
All masking variables are learned as real numbers in $[0,1]$ during training. We follow Louizos et al. (2017); Guo et al. (2023) and employ the augmented $L_{0}$ regularization:
$\mathbf{u}$ $\sim U(0,1)$ (4)
$\mathbf{s}$ $=\operatorname{sigmoid}\left(\frac{1}{\beta}\log\frac{\mathbf{u}}{\mathbf{1-u}}+\log\alpha\right)$
$\tilde{\mathbf{s}}$ $=\mathbf{s}\times(r-l)+l$
$\mathbf{z}$ $=\min(1,\max(0,\tilde{\mathbf{s}})),$
where $\mathbf{u}$ is uniformly sampled between 0 to 1. $\alpha$ is the parameter to be learned and $\beta$ is a hyperparameter. $l,r$ is often $-0.1$ and $1.1$ to ensure most $\mathbf{z}$ are either 00 or $1$ after training.
To prevent models from drastically converging to different sizes, we follow Wang et al. (2019) to use this Lagrangian term:
$\mathcal{L}_{0}=\lambda_{1}\cdot(\hat{s}-t)+\lambda_{2}\cdot(\hat{s}-t)^{2},$ (5)
where $\lambda_{1}$ and $\lambda_{2}$ are both learnable. This loss term $\mathcal{L}_{0}$ will impose $\hat{s}$ to gradually converge to target sparsity $t$.
Updating parameters with LoRA
Considering massive memory usage during full fine-tuning for LLMs, we incorporate lightweight LoRA Hu et al. (2021) modules into LLM weights to update parameters during pruning.
An incorporated module $W^{\prime}$ is consisted of the original weight $W:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ and sequential LoRA weights parallel to $W$:
$W^{\prime}(X)=W(X)+W_{B}\left(W_{A}(X)\right),$ (6)
where $W_{A}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{r},W_{B}:\mathbb{R}^{r}\rightarrow\mathbb{R}^{m}$ and $r\ll m,n$. During training, $W$ is frozen and only $W_{A}$ and $W_{B}$ are learnable.
Efficient distillation
Instead of simultaneously loading two massive models into memory, we propose to incorporate the frozen and intact model $M$ with different lightweight masks and LoRA modules for the teacher and the student. Formally, let $\mathbf{I}=\{\mathbf{z},\mathbf{W}_{A},\mathbf{W}_{B}\}$ denotes the set of all masks and LoRA modules which is highly parameter efficient ($|\mathbf{I}|\ll|\mathbf{M}|$). By incorporating $\mathbf{I}$ into $\mathbf{M}$, we obtain $\mathbf{M}_{\mathbf{I}}$. The objective of knowledge distillation is the KL-divergence Van Erven and Harremos (2014) between teacher’s and student’s output probability distributions $p$:
$\mathcal{L}_{KL}=D_{KL}(p(\mathbf{M}_{\mathbf{I}_{S}},x),p(\mathbf{M}_{\mathbf{I}_{T}},x)),$ (7)
where $x$ denotes training data. $\mathbf{I}_{S}$ and $\mathbf{I}_{T}$ denote the lightweight modules of student and teacher.
Additionally, intermediate layers of a teacher model can serve as effective targets for training a student model Chen et al. (2021). This objective can be formulated as:
$\mathcal{L}_{layer}=\sum_{l}^{L}\operatorname{MSE}(\mathbf{h}_{l}(\mathbf{M}_{\mathbf{I}_{S}},x),\mathbf{h}_{l}(\mathbf{M}_{\mathbf{I}_{T}},x)),$ (8)
where $\mathbf{h}_{l}$ is the hidden embedding of the $l$-th layer. Therefore, the overall objective is:
$\mathcal{L}=\mathcal{L}_{KL}+\alpha_{1}\mathcal{L}_{layer}+\alpha_{2}\mathcal{L}_{0},$ (9)
where $\alpha_{1},\alpha_{2}$ are hyperparameters to control the importance of different loss terms.
3.2 Progressive Knowledge Distillation with Numerous Teachers
Figure 3: Illustration of the sparsity of teacher and student models during pruning. Take the example with the target sparsity $t=50\%$ and sparsity gap $g=10\%$.
All teachers are collected from the snapshot of students as the dotted line illustrated in Figure 2. To narrow the capacity gap between the intact teacher and high sparsity students, we leverage a novel progressive knowledge distillation (PKD) method for pruning. It consists of two stages when pruning a model from 0% sparsity as illustrated in Figure 3.
Before reaching target sparsity
The sparsity of pruned model gradually increase from 0 to $t$. To narrow the sparsity gap, we set a fixed gap value $g$ and make the pruned model $S$ guided by teachers $T$ whose sparsity $\hat{s}(T)$ is approximately $g$ less than $\hat{s}(S)$: $\hat{s}(T)=\hat{s}(S)-g$. These teachers are snapshots of previous students. The original intact model serves as the teacher for student $\hat{s}(S)<g$.
To avoid collecting too many teachers, we only collect teachers with an interval of $i$. Therefore, for any teacher with sparsity $\hat{s}(T)$, it is responsible for guiding a student set within a range of sparsity. We use $\rightarrow$ to denote the relationship in which a teacher distills knowledge to students.
$T\rightarrow\{S|\hat{s}(T)+g<\hat{s}(S)<\hat{s}(T)+g+i\}.$ (10)
And the intact model $\mathbf{M}=T_{0}$ is responsible for the early students whose sparsity is less than $g+i$:
$T_{0}\rightarrow\{S|\hat{s}(S)<g+i\}.$ (11)
After reaching target sparsity
When the pruned model reaches the target sparsity $t$, we proceed to the second stage of PKD. The model undergoes distillation by all preceding teachers, with a reduction of sparsity in the teachers. This gradual process guides the model’s learning trajectory from weaker to stronger knowledge and from easier to more challenging concepts. Throughout this stage, the sparsity of the pruned model $\hat{s}$ remains close to the target sparsity $t$, while the masks $\mathbf{z}$ and LoRA modeuls $\mathbf{W}_{A},\mathbf{W}_{B}$ are continually optimized.
To receive sufficient instruction from the best model (the intact model $\mathbf{M}$), the teacher model is maintained as $\mathbf{M}$ during the final period.
3.3 Post Fine-tuning
After the pruning phase, to obtain better performance, we undergo a post fine-tuning stage following LLM-Pruner Ma et al. (2023). We fix the masks and only fine-tune LoRA modules on the Standford Alpaca Taori et al. (2023) dataset.
4 Experiments
4.1 Experimental Setup
Datasets
To assess the zero-shot ability of LLMs, we perform zero-shot classification tasks on seven commonsense reasoning benchmarks: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), ARC-challenge Clark et al. (2018), and OpenBookQA (OBQA) Mihaylov et al. (2018). We evaluate the general capcability of LLMs on the perplexity metric with WikiText Merity et al. (2016) and PTB Marcus et al. (1993) dataset. Additionally, We report the results on 5-shot MMLU to evaluate the in-context learning ability Hendrycks et al. (2020), which consists of 57 tasks covering STEM, humanities, social science, etc.
Models
We assess the performance of NutePrune on the LLaMA-1 family Thoppilan et al. (2022), comprising LLaMA-7B, LLaMA-13B. Our experiments primarily center on pruning the LLaMA-7B model to provide a comprehensive comparison with previous studies. We primarily evaluate at sparsity levels of 20% and 50%.
Ratio Tune Method BoolQ PIQA
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
[…]
0% LLaMA-7B 73.18 78.35
72.99
67.01
67.45
41.38
42.40
63.25 66.39
20% LLM-Pruner 59.39 75.57
65.34
61.33
59.18
37.12
39.80
56.82 59.01
LoRAPrune 57.98 75.11
65.81
59.90
62.14
34.59
39.98
56.50 -
[…]
67.96
66.69
63.72
38.05
40.00
59.46 63.03
20%   $\checkmark$ WANDA 65.75 74.70
64.52
59.35
60.65
36.26
39.40
57.23 -
LLM-Pruner 69.54 76.44
68.11
65.11
63.43
37.88
40.00
60.07 61.94
LoRAPrune 65.82 79.31
70.00
62.76
65.87
37.69
39.14
60.05 -
LoRAShear 70.17 76.89
68.69
65.83
64.11
38.77
39.97
60.63 -
Compresso 73.64 75.08
64.77
67.72
66.12
37.54
40.40
60.75 62.60
[…]
68.99
65.51
65.49
38.48
40.20
61.15 63.57
NutePrune 74.56 77.04
70.01
65.67
65.78
37.97
39.20
61.46 64.39
[…]
66.75
62.04
58.08
36.77
39.00
58.01 61.72
[…]
66.69
64.56
61.49
37.03
39.20
58.71 63.12
NutePrune 68.99 77.20
67.90
65.04
63.76
37.80
40.20
60.13 63.78
50% LLM-Pruner 52.32 59.63
35.64
53.20
33.50
27.22
33.40
42.13 40.94
LoRAPrune 51.78 56.90
36.76
53.80
33.82
26.93
33.10
41.87 -
[…]
53.03
57.06
45.45
30.03
36.60
50.35 53.14
50% $\checkmark$ WANDA 50.90 57.38
38.12
55.98
42.68
34.20
38.78
45.43 -
LLM-Pruner 61.47 68.82
47.56
55.09
46.46
28.24
35.20
48.98 48.97
LoRAPrune 61.88 71.53
47.86
55.01
45.13
31.62
34.98
49.71 -
LoRAShear 62.12 71.80
48.01
56.29
47.68
32.26
34.61
50.39 -
Compresso 60.09 66.70
39.31
51.93
48.82
27.82
33.40
46.87 47.43
[…]
53.87
57.77
46.59
31.74
35.8
51.13 53.94
NutePrune 62.26 71.00
55.88
57.54
51.68
32.17
34.40
52.13 54.91
  • $\dagger$ only prunes the model by training masks without incorporating LoRA modules.
  • $\ddagger$ prunes the model by co-training the masks and LoRA modules but without post fine-tuning on Alpaca.
  • $\star$ includes our reproduced results with the newer version of lm-evaluation-harness that fixes a LLaMA tokenization issue. See Appendix A for detailed results.
Table 1: Zero-shot performance (%) of the compressed LLaMA models. Bold denotes the best average performance at the same setting.
Baselines
Numerous works delve into pruning techniques for LLMs, including SparseGPT Frantar and Alistarh (2023) and WANDA Sun et al. (2023). Considering the benefits of inference acceleration, we mainly focus on structure pruning methods. For open-source methods as far as we know, we implement LLM-Pruner Ma et al. (2023) and Compresso Guo et al. (2023) and conduct detailed comparisons with our NutePrune. For more recent works that are not publicly available, We assess NutePrune using the same settings as theirs for comparison with their reported results. This includes LoRAPrune Zhang et al. (2023) and LoRAShear Chen et al. (2023).
Implementation details
For pruning stage, we sample 20,000 sentences from the C4 Raffel et al. (2020) dataset with a length of 512 tokens. We train with AdamW optimizer, a batch size of 16, and learning rates of 0.1 for masks and 0.001 for LoRA. We prune the model for 7 epochs. We use a linear sparsity schedule to gradually increase the target sparsity from 0 to the target ratio. We set sparsity warmup to 4 epochs for pruning of 20% sparsity and 1 epoch for 50%. The sparsity gap between the teacher and student $g$ is 10% and the snapshot interval $i$ of teachers is 1%. After pruning, we post fine-tune the pruned model on the Alpaca dataset Taori et al. (2023) for 3 epochs. All experiments are conducted on one A100 GPU (80G).
4.2 Results
Zero-shot performance
Table 1 demonstrates the zero-shot performance on commonsense reasoning tasks for compressed LLaMA models. We mainly present results with the previous version of lm-evaluation-harness for comprehensive comparison, which is widely employed in prior studies Ma et al. (2023). Additionally, we present results with a newer version that addressed a tokenization issue in LLaMA for more accurate evaluation, where we report our reproduced results of publicly available methods. The reported results include experiments for 20%, 25% and 50% sparsity levels, covering scenarios with and without parameter tuning.
The average performance of NutePrune consistently outperforms previous methods across all settings. For pruning without tuning, NutePrune outperforms LLM-Pruner by 2.64%/8.22% at 20%/50% sparsity, underscoring its ability to derive a more effective pruned structure compared to other methods. For pruning with LoRA con-trained, NutePrune improves from 59.46%/50.35% to 61.46%/52.13% at 20%/50% sparsity, indicating co-training with LoRA could help recover model capability damaged by pruning. And with addtional post fine-tuning on Alpaca, notably, it retains 97.17% of the performance of the original model at 20% sparsity and 95.07% at 25% sparsity. NutePrune exhibits more noticeable improvements at higher sparsity levels, proving the effectiveness of our PKD in mitigating the capacity gap between teacher and student. Specifically, NutePrune achieves comparable performance to other methods at 20% sparsity when operating at 25% sparsity. And even without parameter tuning, NutePrune achieves results comparable to other methods at 50% sparsity. These findings suggest that NutePrune is particularly effective in compressing models to higher sparsity levels.
Perplexity
The perplexity (PPL) is a metric reflecting the generation ability of LLMs. The results on WikiText2 and PTB are presented in Table 2. Compresso prunes on an instruction dataset and depends on a meticulously crafted instructing prompt, enabling it to outperform LLM-Pruner on zero-shot tasks at low sparsity but yielding inferior performance on PPL, especially at high sparsity. In contrast, NutePrune utilizes a more general 20k C4 dataset, preserving overall language ability. This distinction allows NutePrune to avoid degradation of generalization and achieve superior performance in terms of PPL compared to both Compresso and LLM-Pruner.
[…]
0% LLaMA-7B 5.68 8.81
20% LLM-Pruner 9.96 15.61
[…]
20% $\checkmark$ LLM-Pruner 8.57 12.84
Compresso 10.38 16.14
NutePrune 7.65 12.40
[…]
25% $\checkmark$ NutePrune 7.85 13.10
50% LLM-Pruner 98.1 224.54
[…]
50% $\checkmark$ LLM-Pruner 22.76 33.9
Compresso 59.73 80.38
NutePrune 13.20 21.97
Table 2: Perplexity metric on WikiText2 and PTB.
Pruning of larger model
In addition to evaluating LLaMA-7B, we assess the larger LLaMA-13B with 20% sparsity. As demonstrated in Table 3, our approach yield an average zero-shot performance of 67.51%, which is only 0.12% lower than the full model and 1.75% higher than LLM-Pruner. It also outperforms LLM-Pruner in terms of PPL. Besides, in-context learning is a crucial capability for LLMs Brown et al. (2020), particularly for larger language models such as LLaMA-13B. To assess it, we conduct evaluations using the MMLU with a 5-shot setting. Results demonstrates that NutePrune can also perform well on MMLU metrics with larger models.
Method Zero-Shot (%) WikiText2$\downarrow$ MMLU
0% Baseline 67.63 5.62 0.426
LLM-Pruner 65.76 6.95 0.351
NutePrune 67.51 6.55 0.355
Table 3: Performance of the compressed LLaMA-13B models with 20% sparsity.
Inference latency
We test the inference latency by generating from 64 tokens to 256 tokens on vLLM Kwon et al. (2023), which is a fast and widely deployed library for LLM inference and serving. The results are presented in Table 4. NutePrune achieves latency savings of 11% and 29% at sparsity levels of 20% and 50%, respectively. While LLM-Pruner can save slightly more latency due to its predefined neater structure, it comes at the cost of reduced flexibility in tailoring. However, as sparsity increases, the difference becomes negligible.
Method 20% 50%
0% Baseline 3.06
LLM-Pruner 2.63(-14%) 2.17(-29%)
NutePrune 2.72(-11%) 2.18(-29%)
Table 4: Inference latency of pruned LLaMA-7B.
Training cost
We report the memory and latency cost on different settings in Table 5. For extra GPU memory cost of PKD, NutePrune snapshot lightweight modules (masks and LoRA) of numerous teachers into CPU. Only one teacher module is loaded onto the GPU when needed, resulting in negligible memory cost compared with KD. In terms of extra time cost, compared with supervised training, KD requires one extra forward pass of teacher model, which is inevitable and cost 18.0% extra latency. When snapshoting a teacher or switching to a new teacher, due to the extremely low frequency of operations, the time can be ignored. Introducing $\mathcal{L}_{layer}$ requires additional 32% memory which is also efficient compared to conventional KD.
Progressive KD $\mathcal{L}_{layer}$ Memory Latency
27.68 3.67
$\checkmark$ 28.67 4.33
$\checkmark$ $\checkmark$ 28.69 4.33
$\checkmark$ $\checkmark$ $\checkmark$ 38.00 5.52
Table 5: Training cost measured by average GPU memory (GB) and per step latency (s/iter).
4.3 Ablation Study
We validate the effectiveness of NutePrune and investigate which properties make for a good NutePrune. Results are average zero-shot performance with tuning but without post fine-tuning, unless otherwise stated.
Effectiveness of PKD
To validate progressive knowledege distillation (PKD) in our NutePrune, we conduct ablation studies on various learning strategies. We eliminate the progressive schedule and adopt standard KD, where the intact model serves as the teacher throughout. Subsequently, we exclude the entire distillation procedure and employ the standard generative language model loss, specifically next-token prediction, to train masks and LoRA modules. The results presented in Table 6 demonstrate the critical role of KD in enhancing performance, with further improvements achieved through PKD. This phenomenon is particularly pronounced at higher sparsity.
Progressive KD 20% 50%
$\checkmark$ $\checkmark$ 63.57 53.94
$\checkmark$ 63.19 52.73
59.98 41.77
Table 6: NutePrune and variants at 20%/50% sparsity.
Two stages of PKD
PKD includes one stage before reaching target sparsity and the other stage after that. Different progressive schedules are adopted. To assess the effectiveness of them, we conducted an ablation study at 50% sparsity under two training settings, as shown in Table 7: training masks only and co-training masks with LoRA. In a stage without a progressive schedule, the intact model serves as the teacher. For the masks-only scenario, adopting either stage 1 or 2 alone yields significant improvements over KD. And for co-training, significant improvement is observed when both stages are adopted simultaneously.
Stage 1 Stage 2 Avg.(%)
masks-only co-train
$\checkmark$ $\checkmark$ 53.14 53.94
$\checkmark$ 52.31 52.79
$\checkmark$ 52.40 52.53
51.83 52.73
Table 7: Performance of two stages of PKD.
Sparsity gap between teacher and student
During stage 1, the sparsity gap between teacher and student model is an important hyperparameter. We conduct experiments with various sparsity gaps and the results are presented in Table 8. A 10% gap is deemed appropriate to prevent a gap that is too small, as it may result in insufficient guidance, or a gap that is too large, which would also toughing distillation.
Sparsity Gap Avg.(%)
5% 53.62
10% (ours) 53.94
20% 53.04
Table 8: Performance of various sparsity gap.
Snapshot interval of teachers
When taking snapshots of students as teachers, it is preferable to save as many teachers as possible to facilitate more comprehensive training. However, the additional CPU storage for these teachers incurs extra costs. As demonstrated in Table 9, selecting an interval of 1% leads to significant improvement over the 10% interval, and the associated extra storage is acceptable, which means we use 10 teachers for 20% sparsity and 40 teachers for 50% sparsity.
Snapshot Interval CPU Storage Avg.(%)
1% (ours) 728MB 53.94
10% 73MB 53.27
Table 9: Storage and performance of various intervals.
5 Conclusion
In this work, we propose NutePrune as a novel efficient pregressive structured pruning method for LLMs. By minimizing the cost of KD, NutePrune aligns closely with the memory and time requirements of conventional training, allowing the utilization of numerous teachers for distillation. This approach mitigates the capacity gap between teacher and student, preserving the quality of the distillation process. We showcase the effectiveness of NutePrune, particularly at high sparsity levels, across various tasks and metrics for LLMs. The findings of this work contribute to the advancement of structured pruning techniques for LLMs, particularly in resource-constrained scenarios.
6 Limitations
We acknowledge the main limitation of this work is that we only evaluate our methods on LLaMA with 7B and 13B due to limited computation resources. And other model families, such as LLaMA-v2, are not included due to the absence of results for comparison from state-of-the-art methods. Recent work Ma et al. (2023); Xia et al. (2023) proves that using extensive data for post-training could substantially enhance the performance, but it comes with a substantial increase in computational costs. We target on pruning on resource-constraint scenarios and leave pruning with extensive data for future work.
Appendix A Zero-shot Performance with Newer Version
lm-evaluation-harness released a new version in June 2023 to assess the zero-shot performance of LLaMA  [^3] https://github.com/EleutherAI/lm-evaluation-harness/pull/531. This update addressed a tokenization bug specific to LLaMA, resulting in higher and more accurate performance results compared to the older version. Despite these improvements, current state-of-the-art reports continue to reference the older version. Consequently, we conducted experiments using both the new and old versions, and the detailed results for the new version are presented in Table 10.
Ratio Tune Method BoolQ PIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
0% LLaMA-7B 75.11 79.16 76.21 69.85 75.29 44.71 44.40 66.39
20% LLM-Pruner 57.49 76.06 69.53 63.93 67.17 38.05 40.80 59.01
[…]
20% $\checkmark$ LLM-Pruner 67.37 77.86 71.47 65.90 69.57 39.59 41.80 61.94
Compresso 73.21 75.90 66.90 68.90 69.99 41.47 41.80 62.60
[…]
NutePrune 75.38 78.02 72.97 67.40 73.82 40.36 42.80 64.39
[…]
NutePrune 74.95 77.75 71.27 67.40 71.25 41.81 42.00 63.78
50% LLM-Pruner 46.48 61.10 36.87 51.78 35.10 27.65 27.60 40.94
[…]
50% $\checkmark$ LLM-Pruner 57.89 69.97 50.06 52.64 49.66 28.58 34.00 48.97
Compresso 61.31 66.32 40.73 52.41 51.18 27.65 32.40 47.43
[…]
NutePrune 67.52 71.60 58.64 60.14 59.72 32.94 33.80 54.91
Table 10: Zero-shot performance of the compressed LLaMA models in the new version of lm-evaluation-harness. Bold denotes the best average performance at the same setting.
Appendix B Pruned Structure
To gain insights into the pruned model, we present a detailed overview of the pruned structure at sparsity levels of 20% and 50%. The original hidden dimension is 4096, with a number of heads set at 32 and an intermediate dimension of 11008. Tables 11 and 12 reveal several observations. Notably, NutePrune tends to avoid pruning the hidden dimension, which aligns with the observation that pruning it may result in significant performance degradation Ma et al. (2023). Regarding heads and intermediate dimensions, NutePrune tends to prune the the last few layers. This observation differs from LLM-Pruner, which asserts the importance of the last layers. Further analysis of this phenomenon is left for future work.
# Hidden Dim 4070
Layer 1 2 3 4 5 6 7 8
# Head 23 22 30 22 29 27 30 28
# Intermediate Dim 5832 7820 9169 9187 8967 9163 9186 9112
Layer 9 10 11 12 13 14 15 16
# Head 30 30 31 30 32 27 30 30
# Intermediate Dim 9261 9165 9303 9695 10005 10258 10417 10564
Layer 17 18 19 20 21 22 23 24
# Head 30 29 26 25 23 21 16 21
# Intermediate Dim 10715 10759 10785 10790 10808 10778 10729 10707
Layer 25 26 27 28 29 30 31 32
# Head 14 15 6 7 11 8 7 9
# Intermediate Dim 10568 10366 10054 9403 8519 7297 6588 5164
Table 11: Detailed structure of compressed 20% LLaMA.
# Hidden Dim 4021
Layer 1 2 3 4 5 6 7 8
# Head 24 21 27 18 26 21 20 20
# Intermediate Dim 3980 6216 7120 6590 5889 5731 5283 4944
Layer 9 10 11 12 13 14 15 16
# Head 25 21 27 26 23 21 25 21
# Intermediate Dim 4879 4563 4590 5040 5469 5832 6215 778
Layer 17 18 19 20 21 22 23 24
# Head 23 23 21 19 16 15 6 13
# Intermediate Dim 7706 8121 8005 8091 8232 7705 6879 6134
Layer 25 26 27 28 29 30 31 32
# Head 7 8 2 5 7 5 3 8
# Intermediate Dim 4912 3762 3107 2527 2357 2314 2931 2622
Table 12: Detailed structure of compressed 50% LLaMA.
Appendix C Generated Examples
We present generated examples from our pruned model using NutePrune at 20% sparsity. We provide examples of three types: without tuning (w/o tune), with tuning but without post-finetuning (w/ tune), and with tuning and post fine-tuning (w/ tune + post FT). The results are displayed in Table 13.
Model
Generated Sentences
w/o Tune
Neural Sequential Model, especially transformers, are used to model the human brain. The model is based on the assumption that the brain is a system of interconnected modules, each of which is responsible for a specific function. The modules are connected by synapses, which are the synaptic connections between neurons. The synapsis are responsible for the transfer of information between the modules. The information is transferred in the form of a signal, which is a combination of the input signal and the output signal. The input signal is the signal that is received by the module, while the output is the output of the module. The output signal is then passed on to the next …
AI can create a logo in seconds . It can also create a logo that is more creative than a human designer. AI can also create logos that are more creative and unique than a human can. AI is also able to create logos in a short time. AI has the ability to create logotypes. AI will be able to create a logo for a company in a short period of time. A logo is a symbol that is used to identify a company. A logo can be used to identify the company. AI also has the ability of creating logos that can be used in different ways. AI creates logos that have …
What’s great about the holiday season ? The holiday spirit, of course! But it’s also a time for giving. And giving is what we do best at the Chamber. We give you the opportunity to give back to your community. We’ve got a great list of charitable organizations that are looking for your help. And we’ve made it easy for you to give. Just click on the link below and you’ll be taken to the website of the charity of your choice. You can make a donation online, or you can mail a check. Either way, you’re giving back to the community. And that’s what the …
w/ Tune
Neural Sequential Model, especially transformers, are used to model the human brain. The model is a neural network that is trained to recognize patterns in data. The network is trained by feeding it data and then testing it on new data. This is a form of machine learning. The neural network is a type of artificial neural network. The term neural network refers to a group of neurons that are connected together. The neurons are connected by synapses. The synapsis are the connections between the neurons. The connections are called weights. The weights are used by the network to determine how much influence each neuron has on the other neurons in the network …
AI can create a logo in seconds , but it can’t design a logo that will work for a company. AI is great for creating a logo, but not for designing a logo. A logo is a visual representation of a company’s brand. It’s a visual symbol that represents the company’ logo. It is a symbol that is used to identify a company and its products. A good logo is one that is memorable and easy to recognize. A great logo is memorably and easy recognizable. A bad logo is not memorable or easy to recognize. A poorly designed logo is also not memorably or easy recognized. A …
What’s great about the holiday season is that it’s a time for giving. And what better way to give than to give the gift of a new home? Whether you’re buying a home for the first time or you’ve been in the market for a while, there are a few things you should know before you start looking. 1. Know your budget. This is the most important thing to consider when you’ll be looking for a home. You’ll want to know how much you can afford to spend on a home, and you’d want to make sure you don’t go over your budget when you start searching for homes …
w/ Tune + Post FT
Neural Sequential Model, especially transformers, is a popular model for deep learning. It is a model that can be trained on large datasets. It is used for tasks such as image classification, natural language processing, and other tasks. It is an effective model for tasks such as image classification and natural language processing. It is also used for tasks like natural language understanding and natural film fiction. It is able to learn from large datasets and can be trained for various tasks. It can be used for tasks that require a large amount of data. It is capable of learning from large datasets. It is used to train neural networks for tasks such image classification …
AI can create a logo in seconds. AI Logo Designer. A Logo AI Can Create. A logo is a key element of any business. It helps to create a unique identity and helps to stand out from the crowd. A good logo can help to build trust and credibility. It can also help to increase the sales. Alogoai.com is a AI logo design service. We can design logos for any business, with the latest technology. We provide high quality logos at affordable prices. Logo design service provided by Alogo AI. Logos designed by AI are more accurate and effective. Logs designed by …
What’s great about the holiday season is the opportunity to give back to the community. Whether it’s donating to a local charity or giving to a needy family, it’ll make you feel good to know that you’re making a difference. Here are some ideas to get started. Volunteer at a local organization. If you have some free time, consider volunteering at a nonprofit organization. You can help with the holidays by helping with the decorations, helping with food preparation, or even helping with some of the administrative tasks. You’ll feel great knowing that you are making a contribution to the organization. Donate …
Table 13: Generated Examples from the Compressed LLaMA-7B at 20% sparsity